Two populations are considered: cell phone users and nonusers. n1 = 98 cell phone users and n2=102 non-users are randomly selected and checked for brain cancer. Is the proportion of brain cancer different in these 2 populations? Perform the Z test for comparison of proportions (Normal Theory method) to test

Cell-users Non-users Total
Brain cancer 18 7
No brain cancer 80 95
Total 98 102

Ho: pbc cell-user =pbc non-cell-useri.e. Ho: p1=p2

Ha: pbc cell-user ≠ pbc non-cell-user Ha: p1 ≠ p2

Find the proportion of non-cell-users who have developed brain cancer.

a.

.0686

b.

The correct answer does not appear as one of the choices.

c.

.875

d.

.035

e.

7
Answer #1

Find the proportion of non-cell-users who have developed brain cancer.

the proportion of non-cell-users who have developed brain cancer = 7/(7+18) = 7/25 = 0.28

So correct choice is "The correct answer does not appear as one of the choices"

Let's use minitab:

Step 1: Click on Stat >>> Basic Statistics >>>2 Proportions...

Step 2: Select Summarized data

Fill the given information

Look the following picture ...

2 Proportions (Test and Confidence Interval) Samples in one column Samples: Subscripts: Samples in different columns First: Second: Summarized data Trials: 98 102 Events: Eirst: 18 Second:7 Select Options... Help OK Cancel

Then click on Option:

Look the following image:

2 Proportions - Options Confidence level: 95.0 Test difference: 0.0 Alternative: notequa 反IJsepooled.estimateofp.fortesti Hel

Then click on OK again click on Ok

So we get the following output

Test and Cl for Two Proportions Sample XN Sample p 2 18 98 0.183673 7 102 0.068627 Difference = p (1) -p (2) Estimate for difference: 0.115046 95% CI for difference : (0.0240264, 0.206066) Test for difference = 0 (vs not = 0): Z = 2.46 P-value 0.014 Fishers exact test: P-Value0.018

From the above output

z = 0.2.46 , p-value = 0.014

Decision rule:

1) If p-value < level of significance (alpha) then we reject null hypothesis

2) If p-value > level of significance (alpha) then we fail to reject null hypothesis.

Here p value = 0.014 < 0.05 so we used first rule.

That is we reject null hypothesis

Conclusion: At 5% level of significance there are sufficient evidence to say that the sample data indicates the proportion of brain cancer different in these 2 populations